# Optimizing Charge-coupled Device Readout Enabled by the Floating-Gate Amplifier

### v1

@@ Abstract
Multiple-Amplifier Sensing (MAS) charge-coupled devices (CCDs) have recently been shown to be promising silicon detectors that meet noise sensitivity requirements for next generation Stage-5 spectroscopic surveys and potentially, future space-based imaging of extremely faint objects such as the Habitable Worlds Observatory. Building upon the capability of the Skipper CCD to achieve deeply sub-electron noise floors, MAS CCDs utilize multiple floating-gate amplifiers along the serial register to increase the readout speed by a factor of the number of output nodes compared to a Skipper CCD. We introduce and experimentally demonstrate on a 16-channel prototype device new readout techniques that exploit the MAS CCD’s floating-gate amplifiers to optimize the correlated double sampling (CDS) by resetting once per line instead of once per pixel. With this new mode, we find an optimal filter to subtract the noise from the signal during read out. We also take advantage of the MAS CCD’s structure to tune the read time by independently changing integration times for the signal and reference level. Together with optimal weighted averaging of the 16 outputs, these approaches enable us to reach a sub-electron noise of 0.9 e- rms pix-1 at 19 $\mu$s pix-1 for a single charge measurement per pixel — simultaneously giving a 30% faster readout time and 10% lower read noise compared to performance previously evaluated without these techniques.
@@ 1 Introduction
Silicon charge-coupled devices (CCDs) have been prized in ground and space-based optical astronomical imaging and spectroscopy for its linearity, low-noise, quantum efficiency, and flight heritage. As the ambitious scientific drivers for planned missions such as a Stage-5 (Spec-S5) spectroscopic survey for precision cosmology or the Habitable Worlds Observatory (HWO) for direct exo-Earth detection demand ever lower noise floors down to the quantum limit, new detector architectures must be developed to meet aggressive sensitivity (e.g., $<1$ e- rms pix-1) and fast readout (e.g., $\lesssim 14$ $\mu$s pix-1) requirements (Rauscher et al., 2016; Asai et al., 2024). Many superconducting quantum detectors have been proposed and tested but still have technical challenges that hinder wider adoption in astronomy, including multiplexing into large pixel array formats, specialized fabrication processes, and demanding cryogenic cooling and infrared radiation shielding requirements that significantly increase engineering complexity (e.g., Berggren et al., 2013; Shaw et al., 2024). Thus, the Skipper CCD has garnered interest as a potential solution because of its photon-counting capability that also retains the favorable properties of modern scientific CCD sensors and only requires passive cooling. Skipper CCDs have already proven to be scientifically valuable in several ongoing sub-GeV direct dark matter search experiments with world-leading constraints (e.g., Abramoff et al., 2019; Arnquist et al., 2023; Adari et al., 2025).
@@
In this paper, we introduce readout techniques enabled by the floating-gate amplifiers of the MAS CCD aimed at further reducing the noise floor with a single read per amplifier while also decreasing the required pixel readout time. A modified sense node reset scheme is implemented, additional optimization in the pixel sampling process downstream from the controller is explored, and a method for optimally averaging amplifier images on the overall noise performance is studied. We experimentally demonstrate these techniques on a prototype 16-channel MAS CCD using Hydra readout electronics, a 16-channel adaptation of the DESI front-end electronics (FEE), both of which are previously described in Lin et al. (2024b). These improvements show the potential of an extended iteration of the MAS CCD with double the number of channels and differential readout to satisfy or even exceed the requirements of future massively multiplexed spectroscopic surveys.
This paper is structured as follows: Section 2 describes the readout approaches and optimal filtering approach we explore in the context of the floating-gate amplifier, Section 3 summarizes our results using the optimal filter to improve the read noise and read time, Section 4 describes additional efforts to improve the read noise by optimally combining the amplifiers, and Section 5 summarizes our main findings and comments on implications for future work. We deploy the techniques discussed in this paper for the readout of a thick, 16-channel MAS CCD with DESI electronics. Details regarding the detector, package, electronics, and baseline performance from experimental tests are found in Lin et al. (2024b), where a read noise of 1.03 e- pix-1 at a read time of 26 $\mu$s pix-1 was reported. The experimental setup used in this paper is similar to that previously described in Lin et al. (2024b), but instead of liquid nitrogen cooling, an integral Stirling cryocooler (Sunpower CryoTel® GT) now provides the cooling power for the 140 K operating temperature of the CCD.
@@ 4 Optimal Amplifier Averaging
We apply a similar prescription outlined in Lapi et al. (2024) for calculating these optimal weights and the approach to the one considered for the optimal filter. We estimate the ideal weights by a least squares minimization where a constraint is now imposed such that the sum of the weights is equal to one, which was not necessary for the filter normalization because it was set by the comparison with the signal level. We adopt an approach using Lagrange multipliers to optimize these weights on each amplifier to minimize the noise of the combined image. When combining the images normally, each amplifier is given an equal weight of one and averaged over the number of amplifiers. To tune these weights $w_{a}$ and recover the optimally weighted average image, we apply the $\sum_{a}^{N_{a}}w_{a}=1$ constraint. As a result, we have the Lagrangian function
@@ 5 Summary
We also found that for most integration times, there was minimal cross-amplifier noise. By averaging the 16 amplifier images using optimal weights and comparing with averaging assuming uncorrelated amplifiers, we found that the noise measurements are mostly consistent. In these cases, the correlations between amplifiers are weak, with all coefficients less than $|0.1|$. With these weak correlations, the optimal averaging, compared to the inverse variance averaging, gives at most a 0.03 e- rms pix-1 reduction in the noise. However, we also find that at certain specific integration times where there is correlated noise, a larger improvement of nearly 0.1 $e^{-}$ may result from optimally averaging the information from the 16 amplifiers. While we expect that the particular features we observed will be specific to the readout electronics used, we have demonstrated that the optimal averaging of multiple measurements can provide a more precise measurement of the pixel charge.

### v2

@@ Abstract
Multiple-Amplifier Sensing (MAS) charge-coupled devices (CCDs) have recently been shown to be promising silicon detectors that meet noise sensitivity requirements for next generation Stage-5 spectroscopic surveys and potentially, future space-based imaging of extremely faint objects on missions such as the Habitable Worlds Observatory. Building upon the capability of the Skipper CCD to achieve deeply sub-electron noise floors, MAS CCDs utilize multiple floating-gate amplifiers along the serial register to increase the readout speed by a factor of the number of output nodes compared to a Skipper CCD. We introduce and experimentally demonstrate on a 16-channel prototype device new readout techniques that exploit the MAS CCD’s floating-gate amplifiers to optimize the correlated double sampling (CDS) by resetting once per line instead of once per pixel. With this new mode, we find an optimal filter to subtract the noise from the signal during read out. We also take advantage of the MAS CCD’s structure to tune the read time by independently changing integration times for the signal and reference level. Together with optimal weighted averaging of the 16 outputs, these approaches enable us to reach a sub-electron noise of 0.9 e- rms pix-1 at 19 $\mu$s pix-1 for a single charge measurement per pixel — simultaneously giving a 30% faster readout time and 10% lower read noise compared to performance previously evaluated without these techniques.
@@ 1 Introduction
Silicon charge-coupled devices (CCDs) have been prized in ground and space-based optical astronomical imaging and spectroscopy for their linearity, low-noise, quantum efficiency, and flight heritage. As the ambitious scientific drivers for planned missions such as a Stage-5 (Spec-S5) spectroscopic survey for precision cosmology or the Habitable Worlds Observatory (HWO) for direct exo-Earth detection demand ever lower noise floors down to the quantum limit, new detector architectures must be developed to meet aggressive sensitivity (e.g., $<1$ e- rms pix-1) and fast readout (e.g., $\lesssim 14$ $\mu$s pix-1) requirements (Besuner et al., 2025; Rauscher et al., 2016; Asai et al., 2024). Many superconducting quantum detectors have been proposed and tested but still have technical challenges that hinder wider adoption in astronomy, including multiplexing into large pixel array formats, specialized fabrication processes, and demanding cryogenic cooling and infrared radiation shielding requirements that significantly increase engineering complexity (e.g., Berggren et al., 2013; Shaw et al., 2024). Thus, the Skipper CCD has garnered interest as a potential solution because of its photon-counting capability that also retains the favorable properties of modern scientific CCD sensors and only requires passive cooling. Skipper CCDs have already proven to be scientifically valuable in several ongoing sub-GeV direct dark matter search experiments with world-leading constraints (e.g., Abramoff et al., 2019; Arnquist et al., 2023; Adari et al., 2025).
@@
In this paper, we introduce readout techniques enabled by the floating-gate amplifiers of the MAS CCD aimed at further reducing the noise floor with a single read per amplifier while also decreasing the required pixel readout time. A modified sense node reset scheme is implemented, additional optimization in the pixel sampling process downstream from the controller is explored, and a method for optimally averaging amplifier images on the overall noise performance is studied. We experimentally demonstrate these techniques on a prototype 16-channel MAS CCD using Hydra readout electronics, a 16-channel adaptation of the DESI front-end electronics (FEE), both of which were previously described in Lin et al. (2024b). These improvements show the potential of an extended iteration of the MAS CCD with double the number of channels and differential readout to satisfy or even exceed the requirements of future massively multiplexed spectroscopic surveys.
This paper is structured as follows: Section 2 describes the readout approaches and optimal filtering approach we explore in the context of the floating-gate amplifier, Section 3 summarizes our results using the optimal filter to improve the read noise and read time, Section 4 describes additional efforts to improve the read noise by optimally combining the amplifiers, and Section 5 summarizes our main findings and comments on implications for future work. We deploy the techniques discussed in this paper for the readout of a thick, 16-channel MAS CCD with DESI electronics. Details regarding the detector, package, electronics, and baseline performance from experimental tests are found in Lin et al. (2024b), where a read noise of 1.03 e- pix-1 at a read time of 26 $\mu$s pix-1 was reported. The experimental setup used in this paper is similar to that in Lin et al. (2024b), but instead of liquid nitrogen cooling, an integral Stirling cryocooler (Sunpower CryoTel® GT) now provides the cooling power for the 140 K operating temperature of the CCD.
@@ 4 Optimal Amplifier Averaging
We apply a similar prescription outlined in Lapi et al. (2024) for calculating these optimal weights and the approach to the one considered for the optimal filter. We estimate the ideal weights by a least squares minimization where a constraint is now imposed such that the sum of the weights is equal to one, which was not necessary for the filter normalization because it was set by the comparison with the signal level. We adopt an approach using Lagrange multipliers to optimize these weights on each amplifier to minimize the noise of the combined image. When combining the images normally, each amplifier is given an equal weight of one and averaged over the number of amplifiers. To tune these weights $w_{a}$ and recover the optimally weighted average image, we apply the $\sum_{a}^{N_{a}}w_{a}=1$ constraint. As a result, we have the Lagrangian function:
@@ 5 Summary
We also found that for most integration times, there was minimal cross-amplifier noise. By averaging the 16 amplifier images using optimal weights and comparing with averaging assuming uncorrelated amplifiers, we found that the noise measurements are mostly consistent. In these cases, the correlations between amplifiers are weak, with all coefficients less than $|0.1|$. With these weak correlations, the optimal averaging, compared to the inverse variance averaging, gives at most a 0.03 e- rms pix-1 reduction in the noise. However, we also find that at certain specific integration times where there is correlated noise, a larger improvement of nearly 0.1 e- may result from optimally averaging the information from the 16 amplifiers. While we expect that the particular features we observed will be specific to the readout electronics used, we have demonstrated that the optimal averaging of multiple measurements can provide a more precise measurement of the pixel charge.